Hello. I know this is kind of a long question, but I am having issues with it. Any help would be appreciated.If the number of new cars built (in thousands) each year from 1991-1995 is approximated by the quadratic function f(x)= 7.1 x^2+11.53x + 13694, where x=0 represents 1991, x=1 represents 1992, and so on, then use the function to answer the following questions. (a) Approximate the number of cars built in 1993. Round your answer to the nearest whole number. (b) If the model is used to project the number of cars built in future years, how many cars were expected to be built in 2001? Round your answer to the nearest whole number.
Answer by rothauserc(4718)

f(x)= 7.1 x^2 + 11.53x + 13694, where x=0 represents 1991
a) f(2) = 7.1(2^2) + 11.53(2) + 13694 = 13745.46 = approx. 13745
b) 2001 - 1991 = 10
f(10) = 7.1(10^2) + 11.53(10) + 13694 = 14519.3 = approx. 14519